Solve
5 over 3 minus 2 over x-4 equals 1 over 2

Answer by stanbon(75887)

Solve
5 over 3 minus 2 over x-4 equals 1 over 2
-----
5/3 - 2/(x-4) = 1/2
------
Multiply thru by 6(x-4)to get:
5*2(x-4) - 2*3*2 = 3(x-4)
------
10x-40 - 12 = 3x-12
---
7x = 40
x = 40/7
=====================
Cheers,
Stan H.
=====================